Shiftsum for Eve. Jan 9 00


Subject: Shiftsum for Eve. Jan 9 00
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:12:46 CST


                Evening Shift 7 Jan 00
 
On Shift: SL,RF

Hours of beam available at all intensities: 8.5
Hours of beam available at nominal intensity: 8.5
Hours of E799 data (special runs not included): 7.0
Hours of special runs (laser scans, muon runs...): 0.3
Accounting of lost time:
  KFRENDs, ETOT LV pover strip trip

========================================================================
Tapes brought to FCC this shift:
                                
========================================================================
Accesses during shift: None

========================================================================
Runs ended this shift:

15447 E799 data 7.9M events
15449 E799 data 4.9M events

 
=======================================================================
Overview of shift: Fairly smooth E799 data. ETOT LV power strip trip

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

15447 1530 Run ongoing as shift starts.
                        
                1540 Noticed large amplitudes on spill scope and that
                        the live time was down. Called MCR to work on it.
                        The spills get even worse as QXR learns in changes
                        from operators.

        152 1630 Live time still bad. request MCR reduce beam intensity
                        to 8.5 as they continue to work on the problem.
                        Trigger hangs just when MCR thinks they have
                        improved the spill structure.

                1655 Red button on ETOT LV - EW Also HCC illum is half
                        black. Paged Jason, who just got home, to come back.

                1730 JH,RF make access to investigate HCC LV problem.
                        Discovered tripped power strip. Reset; all ok.
                        Also SL replaces ADC module in Cr. 11 Slot 11 at
                        Leo's request

15449 1802 Access over, beam back. Start run 15449.

                1920 Spill structure seems to have improved. Request 9e12.
                        See logbook for live time plot.

                2004 "I'm full" Change tapes, NM4AN pos->neg. Loaded new
                        BA HVs. Unfortunately HVs were set to 0 by mistake,
                        so muon run has only pedestals in it.

15450 2018 Start muon run

                2028 Stop muon run

15451 2035 Start BEAM_799 run@ 9e12

                21:00 fixed online displays with FIX_DISP_ALL
                        we need drt_fresh_start next run.

                22:30 Started BA tests: check 11bit ADCs
                        Summary: all modules are fine (no dead channels or bad
                        channels) Full readout without inputs plugged in have
                        nice pedestal distributions. Problems occure after
                        plugging in input cables. Few channels are dead, many
                        have lowed pedestals. Crate 12 has 2 trerminated
                        FERA ADCs. It does not give any problems now but
                        it needs to be replaced (unterminated)
                        (more details are in the logbook)

                2255 Raised intensity request to 9.5e12 to see if live
                        time has improved. (a little but still low)

========================================================================
Items to Follow up on Next Shift
  Next run do DRT_FRESH _START
  
  Label 30 QKE tapes for short reg run and other special runs
 
  Don't forget to hit "reset quartz" button after each KFREND.



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:12:52 CST